HUBBLE SPACE TELESCOPE DAILY REPORT #2820 PERIOD COVERED: 0000Z (UTC) 03/01/01 - 0000Z (UTC) 03/02/01 Daily Status Report as of 061/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD/MA2 8770 (The Nucleosynthesis of Boron - Benchmarks for the Galactic Disk) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to determine boron (B) abundance. In spite of several HST cycles of B abundance determinations, we do not know the true, ambient, present-day abundance of B. Only two stars {with undepleted Li and Be} have been observed to derive the Population I true B abundance. Past cycles have focused on {1} halo stars with low metallicity and {2} solar-type stars with Li and/or Be deficiencies. In this proposal, the stellar abundances of B will be determined, measured from the B I 2497 Angstrom line from STIS echelle spectra, of main sequence F-G stars that have retained their full initial abundances of boron. Our target stars are those for which ground-based observations show that beryllium is undepleted. These new abundances will map evolution of the B abundance in the Galactic disk in the metallicity range from one-third solar up to solar. Then the B, and previously determined Li and Be abundances, will be used to improve understanding of the nucleosynthesis of B. There were no reported anomalies. 1.2 Completed Two Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.3 Completed Four Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The proposal completed with no reported problems. 1.4 Completed Three Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.5 Completed Two Sets of WF/PC-2 8583 (Imaging Snapshots of Asteroids) The WF/PC-2 was used to obtain images of the fifty largest main belt asteroids that have favorable apparitions during cycle 9. The images will be searched for companion bodies, as well as mineralogical variegation on the resolved main bodies. There were no reported problems. 1.6 Completed Three Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution. The proposal completed nominally. 1.7 Completed FGS-1 8731 (A High Angular Resolution Survey of the Most Massive Stars in the SMC) Fine Guidance Sensor #1 was used to conduct high angular resolution observations of some of the SMC's hottest and most luminous objects by conducting a comprehensive survey of nearly every spectral type in the upper portion of the HR diagram. Binary or multiple star systems will be detected down to an unprecedented level, more than one order of magnitude better than possible with WFPC2. The targets selected include a representative list of normal Main Sequence O-stars and their evolved descendents, namely supergiants, hypergiants, LBVs, and WRs, many of which have been previously observed by HST's spectrometers for purposes of cosmological calibrations. The observations completed, and no problems were reported. 1.8 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.9 Completed FGS/1 8616 (Masses of Pre-Main Sequence Binaries) The Fine Guidance Sensor #1 was used to continue mapping the orbits of young star binaries in the Taurus and Ophiuchus star forming regions. Our goal is to measure their masses dynamically. This is important because there are still no low mass young stars with reliably known masses so calculations of their evolution to the main sequence are uncalibrated. There were no reported problems. 1.10 Completed FGS-1 8783 (Orbits of Pre-Main Sequence Binaries) FGS #1 was used to dynamically measure the masses of low mass pre-main sequence stars. The proposal completed with no reported problems. 1.11 Completed STIS/CCD 8884 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to take parallel images of random areas of the sky. The proposal completed nominally. 1.12 Completed Five Sets of WF/PC-2 8699 (The Origin of Short-Period Comets) The WF/PC-2 was used to detect and characterize cometary nuclei in order to determine the basic physical properties of a large fraction of the population of short-period comets. By acquiring statistically significant data, we can study the origin of this family of comets and test the hypothesis that they are collisional fragments from the Kuiper Belt Objects. The observations completed with no reported problems. 1.13 Completed Two Sets of STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.14 Completed STIS/MA1 8862 (FUV-MAMA Cycle 9 Flats) The Space Telescope Imaging Spectrograph (MA1) was used to obtain FUV-MAMA observations of the internal Krypton lamp to construct an FUV flat applicable to all FUV modes. No problems were encountered. 1.15 Completed STIS/MA2 8863 (NUV-MAMA Cycle 9 Flats) The Space Telescope Imaging Spectrograph (MA2) was used to obtain NUV-MAMA observations of the internal Deuterium lamp to construct an NUV flat applicable to all NUV modes. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 13 Successful: 13 Scheduled Re-acquisitions: 3 Successful: 3 2.2 FHST Updates: Scheduled: 25 Successful: 23 As documented in HSTAR 8092, the full maneuver updates scheduled for 061/002846Z and at 061/003131Z failed due to tracker #3. The subsequent guide star acquisition was successful. 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1A. The engineering status buffer limits were cleared at 061/0848Z, using ROP DF-18A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW